Shift Summary 11/22/99

Richard Ford (rickford@fnal.gov)
Tue, 23 Nov 1999 00:14:41 -0600 (CST)

Shift Summary for eve shift, 22/Nov/99

On Shift: Rick Ford, Julie

Hours of beam available at all intensities: 4.15
Hours of beam available at nominal intensity: 4.15
Hours of E799 data (special runs not included): 3.90
Hours of special runs (laser scans, muon runs...): 0.75

Accounting of lost time: NM1U and NM1H magnet trips due to NS1LCW problems;
tape change.

========================================================================
Runs ended this shift:
Run 15030 - 24.8M events - BEAM_799
Run 15032 - 5.9M events - BEAM_799
Run 15033 - LASER_SCAN

Tapes brought to FCC this shift:
(including tapes for online split)
UP*172-7,UPS058-59

========================================================================
Accesses during shift:
None

========================================================================
Overview of shift:
Run 15032 begins with intermittent missed spills due to magnet/LCW problems.
Beam disappears at 8pm for machine studies.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15030 15:30 Eve shift takes over long E799 data run.

531 16:28 Run ends with tapes full.

Forgot to update tape #s on init.
Fresh start.

15032 16:46 Start run. KQ hangs trigger. Reload maps
and KFREND.

Load UPLA08 on ktev3.

3-43 16:50-17:40 NM1U and NM1H magnets in KTeV beam line
trip off due to problems with NS1LCW
tripping. Expert finally arrives and
fixes it.

18:30 Load UPF046 on ktev2.

18:30 ktev4e and ktev1e 91% full. Call Arun
to have him clean SPLIT areas.

98,103,122,140 18:55-19:50 Live time drops to ~0.40. Spill structure
is notch (see Logbook XXII p.264). SEM
info and livetime recorded (p.283).

149 20:00 Beam disappears. Machine studies begin.

20:05 Push in beam stop crash switch. RickF
leaves. Drop to 1 person shifting.

15033 20:15 Init long laser scan. Try to put in fake
spill signal. Not clear that it works.
Start laser scan anyway.

21:00 Laser scan ends.

15034 21:25 Start long cosmic run. No disk/tapes.

21:30 Turn off NM4AN using AN799POSTONEG. Run
manager will decide whether we need to
polarity.

21:40 Page JasonL for help with the fake spill
signal since it is obviously not working
because DC histos are not updating. He
walks me through the process, since it is
NOT fully documented in the log book.
Fake spill started. DC histos are up.

21:50 RickF arrives. JAWS heads home.

23:58 Night shift arrives. Tschuss.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift